OpenStudy (anonymous):

what are my points on the graph for x+3y=12

jimthompson5910 (jim_thompson5910):

if x = 0, then x+3y = 12 0+3y = 12 3y = 12 y = 12/3 y = 4 giving us one point (0,4)

jimthompson5910 (jim_thompson5910):

to find another point, pick another value of x, plug it in, and solve for y

OpenStudy (rmrjr22):

yup